# Conformal compactification and affine-null metric formulation of the Einstein equations

### v1

@@ IV Summary and discussion
In [60], a partially regularized set of equations in physical space was used to show the echoing of the critical phenomena for the redshift and the corresponding decay of the Bondi mass. The reason for using a partially regularized set of equation was that the compactified Einstein equations in physical space, corresponding to (61), in principle only require a regularization for the metric fields. This is because the compactified evolution equation, (61e), can be easily regularized by moving the singular term at $\hat{x}=1$ from the left hand side to the right hand side. As a result, the time derivative of the scalar field at ’compacified infinity’, $\hat{x}=1$, is zero. The initial scalar field also vanishes on $\hat{x}=1$, due to the Sommerfeld radiation condition (11). The scalar field is therefore always zero on $\mathcal{I}$. Consequently, the monopole can only be numerically determined by calculation of a derivative at $\mathcal{I}$.
The news function driving the Bondi mass decay was not extracted in [60]. By definition the scalar news function is the $u$ derivative of the scalar monopole, thus its numerical calculation on the compactified grid requires to determine a second order mixed $u\hat{x}$-derivative of the scalar field. This numerical derivative was noisy in [60], due to the employed spectral method and the e-folding [49] of the scalar field near the critical solution.
A ‘proper’ scalar field renormalization (where the singular behaviour is absorbed into a new variable, like (64), has the advantage that the news function can easily be extracted by evalulating the right hand side of the compactified and renormalized evolution equations at the ‘location of infinity’, i.e. at $\hat{x}=1$. This avoids calculating derivatives of the scalar field numerically, which can lead to amplification of numerical noise.
Figure 1: Evolution of the initial data profile ${\Phi=\sqrt{12}\Big{[}\arcsin(\frac{b\lambda}{a+\lambda})-\arcsin b\Big{]}}$, with $a=2$ and $b\simeq 0.9472018$. In all three panels, quantities are evaluated at null infinity and the abscissa is the Bondi time $u_{b}$. The upper panel A shows the evolution of the redshift factor $H$ (red dashed line) and the Bondi mass $m_{B}$ (solid black line). The middle panel B shows the news function $N$ (dashed red line) and the scalar monopole $C$ (solid black line). The lower panel C shows the numerical error of the calculation of the right hand side of (51). The vertical dotted lines correspond to the Bondi time of the last null cone connecting with the central geodesic. The horizontal dotted line in panel A and the blue star point out the final value of the Bondi mass at the threshold of black hole formation.
As an example (by adaptation of the code of [60]), we show in panel B of Fig. 1 the news function of a scalar field that evolves to form a black hole. To our knowledge, this is the first extraction of a news function in the context of critical phenomena while using affine-null coordinates, up until now scalar field news functions have been extracted using Bondi-Sachs coordinates by [51, 69, 70]. As explained in Sec. II, the formation of a black hole is indicated if the redshift $H\rightarrow 0$. In panel A of Fig. 1, it is seen that the redshift (red dashed line) drops to zero at the Bondi time $u_{H}:=u_{b}\sim 26.70$. This $u_{b}=const$ slice corresponds to the last null cone that connects an asymptotic observer at null infinity with the central geodesic, where $\hat{x}=0$ and from where the system (35) is integrated towards $\hat{x}=1$. Beyond the time $u_{H}$, the asymptotic observer has no information on how the scalar field continues to collapse toward the central geodesic.
Figure 2: Evolution of the scalar field and its $u-$derivative at the origin in the compactified coordinate where $\lambda=0$. The initial data profile for $\lambda\geq 0$ is the one specified in Fig. 1. The dotted line corresponds to the code time of the last null cone connecting to $\mathcal{I}$.
Panel A also displays the (exponential) decay of the Bondi mass with a final value (blue star) of $m_{B}=1.97\cdot 10^{-2}$ measured at $\mathcal{I}$. The news function (red dashed line) and the scalar monopole (black solid line) are shown in panel B of Fig. 1. In the behavior of the two functions one can see that the news function is the derivative of the monopole with respect to the Bondi time $u_{b}$. It can be understood by correlating the zeros of the news function with the extrema of the scalar monopole. We observe in panel B that they coincide. The Bondi mass decay is coupled with the square of the news function via (51). In panel B of Fig. 1 the right hand side of (51) is shown using the data of the Bondi mass displayed in panel A and the news function in panel B. This balance law was also shown graphically by displaying its components by Barreto [69] in physical space. We can see that the balance law holds up to numerical precision. For about 80% of the overall simulation time of $u_{H}$, the absolute error in the balance law is $\sim 10^{-7}$. It then increases to about $10^{-3}$ until the simulation crashes due to black hole formation. The increase of error can be understood by looking at the evolution of the scalar field at the origin, i.e along the central geodesic. Thereat are the start points of the null rays emanating to null infinity. The scalar field (black curve) and its $u$-derivative (red dashed curve) both evaluated at the origin are shown in Fig. 2. The abscissa is the proper time at the central geodesic and it is also the time measured in the numerical simulation. The central time determines the Bondi time in Fig. 1 by integrating (50). We observe that the zero crossings of the scalar field decrease exponentially towards the final time. The central time $u=6.13$ has the Bondi time $u_{H}$ at $\mathcal{I}$, which is the time when the last null cone emanating from the central is connected with null infinity. The exponential decrease in the zero crossing of the scalar field towards final time, also means that the $u$-derivative of the scalar field along the central geodesic exponentially in amplitude towards the final time (see the red curve in Fig. 2). As the $u$-derivative of the scalar field along the geodesic is determined by a $\hat{x}-$ derivative evaluated at $\hat{x}=0$, as seen in (30) and (34), for times close to black hole formation time, the spatial derivatives at the vertex also become stiffer with increasing amplitudes. As consequence high numerical resolution is required at the vertex to capture the scalar fields dynamical behaviour along the geodesic. It is this stiffness and because of the employed spectral methods in the code of [60], that leads to the increase of error towards $10^{-3}$ in panel C of Fig. 1 and spoils the numerical accuracy in the evaluation of the balance law at times before black hole formation.
@@ Acknowledgements.
E.G. acknowledges financial support from CONICET and SeCyT-UNC. T.M. thanks the organizers of the “Theoretical High-Energy Physics Workshop on Gravity and Holography” in Valparaiso/Viña de Mar, Chile where some of this work was presented.

### v2

@@ V Summary and discussion
In [60], a partially regularized set of equations in physical space was used to show the echoing of the critical phenomena for the redshift and the corresponding decay of the Bondi mass. The reason for using a partially regularized set of equation was that the compactified Einstein equations in physical space, corresponding to (61), in principle only require a regularization for the metric fields. This is because the compactified evolution equation, (61e), can be easily regularized by moving the singular term at $\hat{x}=1$ from the left hand side to the right hand side. As a result, the time derivative of the scalar field at ’compacified infinity’, $\hat{x}=1$, is zero. The initial scalar field also vanishes on $\hat{x}=1$, due to the Sommerfeld radiation condition (11). The scalar field is therefore always zero on $\mathcal{I}$. Consequently, the monopole can only be numerically determined by calculation of a derivative at $\mathcal{I}$. The news function responsible for the Bondi mass decay was not extracted in [60], in Sec.IV we show how this issue can be resolved through a modification of its scheme. The key difference in the new approach is that the regularized scalar field (64) is incorporated into the existing code, thereby enabling a more accurate extraction of the news function; see Fig.2.
@@ Acknowledgements.
E.G. acknowledges financial support from CONICET and SeCyT-UNC. T.M. thanks the organizers of the “Theoretical High-Energy Physics Workshop on Gravity and Holography” in Valparaiso/Viña de Mar, Chile where some of this work was presented. E.G thanks Amato for enlightening discussions.